OpenStudy (anonymous):

Of 13 possible books, you plan to take 8 with you on vacation. How many different collections of 8 books can you take?

OpenStudy (raden):

13C8

OpenStudy (anonymous):

so u just times 8 by 13

OpenStudy (anonymous):

or is it 8!

OpenStudy (raden):

do u know about combination ? and factorial ?

OpenStudy (anonymous):

no

OpenStudy (raden):

hmmmm... for basic, u must know about factorial, first. ex. 3! (3 factorial) it means that 3! = 3 * 2 * 1 = 6 and 8! = 8.7.6.5.4.3.2.1 = ... (use calculator :)) now, the formula of combination is nCr = n!/(n-r)!r!, so 13C8 = 13!/(13-8)!8! = 13!/5!8! = ... calculate it

OpenStudy (anonymous):

51,891,840

OpenStudy (anonymous):

is that right?

OpenStudy (raden):

13!/5!8! =13.12.11.10.9.8!/(5.4.3.2.1.8!) = 13(11)(9) = 1287

OpenStudy (anonymous):

are we using the formula for Permutations?

OpenStudy (raden):

it's combination not permutation

OpenStudy (anonymous):

oh yeah....didn't see that.

OpenStudy (anonymous):

how did u get 11 and 9